A ball is thrown horizontally with a velocity of 12m/s. How long will it take for the ball to hit the ground if it is released 3.5m from the ground.

Assume no air resistance
1
Expert's answer
2020-04-08T10:32:05-0400

The time of motion of a ball

"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2\times 3.5}{9.8}}=0.845\:\rm s."